OpenStudy (anonymous):

Another one I need checked x- 1/8 = 3/4 7/8 -7/8 5/8 -5/8 i picked 5/8 Another one I need checked x- 1/8 = 3/4 7/8 -7/8 5/8 -5/8 i picked 5/8 @Mathematics

OpenStudy (mathmagician):

7/8

OpenStudy (anonymous):

x- 1/8 = 3/4 add 1/8 to both sides x = 3/4 + 1/8 x = 6/8 + 1/8 x = 7/8